Please answer all If you can't answer all then please don't answer just one question. i need all

2. You would like to buy a house in in 16 years and estimate that you will need a deposit of $73,014. You plan to make bi-weekly deposits into an account that you hope will earn 7.05%. How much do you have to deposit every two weeks?
3. You have accumulated $1,085.55 in debt by buying things on Amazon during quarantine.  The minimum monthly payment is $24.87, If you make only the minimum payment, it will take you 65 months to pay it off. If you pay $10 more per month, you will pay it off in 36 months.  How much total interest will you save?
4. Your best friend opened an investment account with a deposit of $185 today. They will continue to deposits every two weeks into an account that they hope will earn an interest rate of 8.48% per year. How much will they have in 5 years?
5. Your grandmother invested in an annuity to help fund her retirement.  She will receive monthly payments of  $4,532 over 18 years and she will receive her first payment today. If the appropriate discount rate is 4.61%, how much is the the annuity worth today?
Answer #1

1.
=PMT(7.05%/26,26*16,0,-73014)=94.9655540955647

2.
=24.87*65-(24.87+10)*36=361.23

3.
=FV(8.48%/26,26*5,-185,0,1)=29990.266769816

4.
=PV(4.61%/12,12*18,-4532)=664369.050392142